calculate the quantity of a gas, in moles, if 45L at 12 celsuis has a pressure of 966mm Hg. the ideal gas constant is 0.0821L atm/mol K
1
Expert's answer
2017-04-05T10:26:07-0400
Solution: 1) Convert data to identical units. Temperature: 12 Celsius = 285.15 Kelvin. Pressure: 966 mm Hg = 1.27 atm.
2) Use ideal gas law to find quantity of a gas. Ideal gas law written as PV = nRT, where P is the pressure of the gas (in atm), V is the volume of the gas (in L), n is the amount of the gas (in moles), R is the ideal gas constant and T is the absolute temperature of the gas (in K). From this equation find quantity of gas by equation: n = PV/RT.
3) Solve the equation. n=(1.27*45*atm*L)/(0.0821*285.15*L*atm*〖mol〗^(-1) )=57.15/23.41=2.44 moles.
Answer: The quantity of the gas is 2.44 moles.